OpenStudy (anonymous):

3x=4y-12 determine x and y-intercept

OpenStudy (anonymous):

3x=4y-12 3x-4y=-12 (ax+by=c) to get x-intercept, substitute y as 0 3x-4(0)=-12 3x=-12 x=-4 to get y-intercept, substitute x as 0 3(0)-4y=-12 -4y=-12 y=3 x-intercept is -4 y-intercept is 3